OpenStudy (anonymous):

once again, need help! find all degree solutions! cos^2 x + 3sin x + 3 = 0

OpenStudy (anonymous):

cos^2 x = 1 - sin^2 x. Does this help at all? :P

OpenStudy (anonymous):

You can rearrange the function to be Sin^2 x - 3sin x - 4 = 0

OpenStudy (anonymous):

Then let sin x = y and solve for y with quadratics

OpenStudy (anonymous):

as in y^2 - 3y - 4 = 0

OpenStudy (anonymous):

i got this far... cos^2 x + 3 sin x +3 = 0 cos^2 x + 3 sin x = -3 cos^2 x = -3 and 3 sin x = - 3 then i did the following: square rooted the cos^2 x so that it became cos x = - sqrt3 for the other is sin x = -1

OpenStudy (anonymous):

cos^2 x + 3 sin x +3 = 0 cos^2 x + 3 sin x = -3 cos^2 x = -3 and 3 sin x = - 3 this part of your thinking is incorrect, because there is a + with those cos and sin. You can't do that, sorryyy :P You have to make all of the trig functions sin or all of them cos, and its way easier to make them all sine, like i showed above

OpenStudy (anonymous):

make cos^2 x into (1-sin^2 x) and then solve for sin^x

OpenStudy (anonymous):

dang it, my brain keeps failing :( i feel like a scrub lol

OpenStudy (anonymous):

still need help, i did the y^2 -3y -4 = 0 and got (y-3)(y-1), but idk what to do from here

OpenStudy (anonymous):

well thats really good. but your brain is failing you. If i will expand what you wrote, i will get y^2 -3y - y + 3 = y^2 -4y + 3

OpenStudy (anonymous):

You mixed your terms up, do you want to try again, or do you just want to know the answer?

OpenStudy (anonymous):

know the answer, i got 4 more to do.... lol

OpenStudy (anonymous):

aight y^2 -3y - 4 = 0 (y+1)(y-4) = 0 y = -1 or 4

OpenStudy (anonymous):

damn it, its (y-3)(y+1) isn't it?

OpenStudy (anonymous):

sin x = -1 or sin x = 4 but sin x can only be between -1 and 1 if you think about it -1 <= sin x <= 1 if you look at sin graph, so the second option is not valid

OpenStudy (anonymous):

Do you follow me here? :D

OpenStudy (anonymous):

 thats definitely how sin curve looks hhahahaah. well it sure doesnt go higher than 1

OpenStudy (anonymous):

so, to find all angles, i must make an equation with sin x = -1, sooooo that means i must find where sin x = -1, which is 270 degrees so my answer is 270 degrees + 360n (n is all real numbers)

OpenStudy (anonymous):

right? riiight? please don't fail me now!

OpenStudy (anonymous):

Yeah the 270 is the only value between 0 and 360 and if you need it forever and in degrees then you are correct

OpenStudy (mathmale):

@sauceboss72: Status? Are you OK with the solution of this problem, or still need help?

OpenStudy (anonymous):

just got done with it, will close here in a minute.

OpenStudy (mathmale):

:)